# Bi-AQUA: Bilateral Control-Based Imitation Learning for Underwater Robot Arms via Lighting-Aware Action Chunking with Transformers

### v1

@@ IV-C2 Generalization to Novel Objects and Disturbances
To evaluate whether Bi-AQUA overfits to the polyurethane block used during training, we assess three additional evaluation settings: a black rubber block, a blue sponge, and bubbles surrounding the manipulated object. The results in Table III demonstrate that Bi-AQUA generalizes well beyond the training distribution despite having only ten demonstrations.
For the black block, Bi-AQUA achieves success rates between 60% and 100%, with perfect performance under white, rgbw, cyan, and purple lighting. This indicates robustness to drastic changes in albedo and object contrast.
For the blue sponge, which introduces significant geometric, textural, and chromatic variation, performance varies more widely (0–100%). Even under these severe appearance shifts, Bi-AQUA maintains non-trivial success rates across multiple lighting modes, achieving 80% under cyan and 40% in the dynamic changing mode.
The bubble condition is particularly challenging from a perception standpoint. Because the top camera observes the scene through the water surface from above, air injection generates surface waves that induce strong refraction and temporal warping, severely degrading the global view of the tank. At the same time, bubbles rise around the gripper and the object, intermittently occluding the local workspace and creating high-frequency specular highlights and false edges in the gripper view. Despite these compounded distortions, Bi-AQUA still achieves up to 100% success in several lighting conditions, as shown in Table III. This indicates substantial resilience to simultaneous global (water-surface) and local (near-gripper) disturbances.
Overall, these results show that Bi-AQUA not only provides strong lighting robustness but also extends to unseen objects and disturbance conditions without any additional data or fine-tuning. The remaining failure cases point to the inherent difficulty of underwater visual degradation and motivate future work on improved relighting models, physics-based augmentation, and multimodal sensing.
@@ V Conclusion
We introduced Bi-AQUA, a bilateral control–based imitation learning framework for underwater manipulation that explicitly models lighting at multiple levels. Through a label-free Lighting Encoder implicitly supervised by the imitation objective, FiLM-based conditioning of visual features, and a sequence-level lighting token added to the transformer encoder input, Bi-AQUA adapts its visuomotor behavior to changing lighting while preserving the force-sensitive advantages of bilateral teleoperation. Ablation and generalization experiments show that the lighting-aware components provide complementary gains and that the learned lighting representation transfers to novel objects and visual disturbances.
Limitations and future work. Our evaluation is restricted to a single task, robot platform, and tank environment with a finite set of lighting modes, and performance still degrades under extreme combinations of lighting and object appearance. We also do not yet consider variation in water quality, background clutter, or large-scale field deployments. Extending Bi-AQUA to richer skill repertoires, more realistic subsea settings, and integration with other adaptation mechanisms (e.g., domain randomization or online adaptation) are important directions for future work.

### v2

@@ IV-D2 Pick-and-Place task: Generalization to Novel Objects and Disturbances
To evaluate whether Bi-AQUA overfits to the polyurethane block used during training, we test three additional settings: a black rubber block, a blue sponge, and bubbles surrounding the manipulated object (Fig. 9). Table III shows that Bi-AQUA generalizes beyond the training distribution, maintaining high success on the black block (60–100%) and non-trivial performance under severe appearance shifts (blue sponge) and compounded visual distortions (bubbles). These results indicate robustness not only to lighting variation but also to object and environmental changes without additional training.
IV-D3 Drawer closing task: Long-Horizon drawer Closing
In the multi-stage drawer closing task, Bi-AQUA consistently outperforms the force-ablated variant across both seen and unseen lighting conditions. Under trained lighting modes, the full model achieves success rates of 100% in red, blue, and green illumination, and 80% in white and rgbw. In contrast, the force-ablated policy exhibits substantial degradation, with success rates dropping to 40–60% in most trained conditions.
Under untrained lighting, the performance gap remains evident. Bi-AQUA maintains high robustness, achieving 100% in cyan and 80% in purple and changing illumination, whereas the force-ablated policy shows inconsistent behavior (40–100%). These results indicate that force integration significantly improves stability in long-horizon sequential manipulation, where perception drift and contact errors may accumulate over multiple stages.
IV-D4 Peg extraction task: Contact-Rich Peg-out-Hole
The performance difference becomes more pronounced in the contact-rich peg extraction task. Without force input, the policy frequently fails, achieving 0% success in red, purple, and changing illumination, and at most 80% even under favorable lighting. In contrast, the full Bi-AQUA model achieves 100% success in all trained lighting modes and maintains high performance under unseen lighting (100% in cyan and changing, 80% in purple).
The large margin in this task highlights the importance of force-aware action chunking in scenarios with tight geometric tolerance and sustained frictional contact. Visual-only control is insufficient when precise alignment and stable interaction are required, particularly under photometric shifts that degrade perception quality.
IV-D5 Summary
Overall, the results demonstrate that Bi-AQUA generalizes beyond the original pick-and-place benchmark to both long-horizon sequential manipulation and contact-rich precision tasks in underwater.
@@ V Conclusion
We proposed Bi-AQUA, a bilateral control-based imitation learning framework for underwater robot arms that explicitly models lighting variation within the visuomotor policy. By combining a label-free Lighting Encoder, FiLM-based visual modulation, and a lighting token for transformer conditioning, Bi-AQUA adapts to challenging underwater illumination while preserving the force-sensitive advantages of bilateral control. Experiments on real underwater manipulation tasks showed that Bi-AQUA outperforms a bilateral baseline without lighting modeling and remains robust under static, unseen, and dynamically changing lighting conditions. These results highlight the importance of explicit lighting-aware adaptation for reliable underwater visuomotor learning and suggest a promising path toward practical autonomous underwater manipulation.